Hometown Bakery sells three types of doughnuts: glazed, jelly, and cake. The following table shows the sales price and variable costs for each type. The bakery incurs $300,000 a year in fixed expenses. Assume that it sells 2 glazed doughnuts for every 1 jelly doughnut and every 1 cake doughnut.

DOUGHNUT TYPE SALES PRICE VARIABLE COST
Glazed $0.35 $0.2
Jelly $0.5 $0.45
Cake $0.40 $0.27

How many doughnuts of each type will be sold at the breakeven point?

Jelly doughnuts

Cake doughnuts

Glazed doughnuts

What amount of revenue would need to be generated by each type of doughnut for the company to earn $60,000 in operating income?

Jelly doughnuts$

Cake doughnuts$

Glazed doughnuts$
Answer #1

1) Weighted average contribution margin per unit = (0.35-.20)*50%+(0.50-0.45*25%)+(0.40-0.27)*25% = 0.12

Break even point = 300000/.12 = 2500000

Jelly doughnuts (2500000*25%) 625000
Cake doughnuts 625000
Glazed doughnuts 1250000

2) required unit = (300000+60000)/0.12 = 3000000

Jelly doughnuts 750000
Cake doughnuts 750000
Glazed doughnuts 1500000